# Neural Dueling Bandits

### v1

@@ Appendix C Theoretical Analysis for Neural Contextual Bandits with Binary Feedback (Section 5)
In this section, we show the proof of 5 and 6 for neural contextual bandits with binary feedback (Section 5). We can largely reuse the proof from Appendix A, and here we will only highlight the changes we need to make to the proof in Appendix A.
@@
To begin with, in our analysis here, we adopt the same requirement on the width of the NN specified in Eq. 10. First of all, 1 still holds in this case, which allows us to approximate the unknown utility function $f$ using a linear function. It is also easy to verify that 2 still holds. As a consequence, 3 and 4 both hold naturally.
C.1 Proof of Confidence Ellipsoid
@@
in which the inequality follows from 3. Then, the proof in Eq. 19 can be reused to show that
@@
Next, making use of the $1$-sub-sub-Gaussianity of the sequence of noise $\{\varepsilon_{s}\}$ and Theorem 1 from [18], we can show that with probability of at least $1-\delta$,
@@
Denote $\varphi(x)=g(x;\theta_{0})$. Recall that 1 tells us that $f(x)=\langle g(x;\theta_{0}),\theta_{f}-\theta_{0}\rangle=\langle\varphi(x),\theta_{f}-\theta_{0}\rangle$ for all $x\in\mathcal{X}_{t},t\in[T]$. To begin with, for all $x\in\mathcal{X}_{t},t\in[T]$ we have that
@@
in which we have used 6 (reproduced in Eq. 38) in the last inequality. Now making use of the equation above and 4, we have that
@@
in which the last inequality follows from Eq. 39 and 4.
@@ C.2 Regret Analysis
Appendix D Leftover details from Section 6
To demonstrate the different performance aspects of our proposed algorithms, we have used different synthetic rewards functions, mainly, $f(x)=10(x^{\top}\theta)^{2}$ (Square) and $f(x)=\cos(3x^{\top}\theta)$. The details of our experiments are as follows: We use a $d$-dimensional space to generate the sample features of each context-arm pair. We denote the context-arm feature vector for context $c_{t}$ and arm $a$ by $x_{t}^{a}$, where $x_{t}^{a}=\left(x_{t,1}^{a},\ldots,x_{t,d}^{a}\right),~{}\forall~{}t\geq 1$. The value of $i$-the element of $x_{t}^{a}$ vector is sampled uniformly at random from $\left(-1,1\right)$. Note that the number of arms remains the same across the rounds, i.e., $K$. We then select a $d$-dimensional vector $\theta$ by sampling uniformly at random from $(-1,1)^{d}$. In all our experiments, the binary preference feedback about $x_{1}$ being preferred over $x_{2}$ (or binary feedback in Section 5) is sampled from a Bernoulli distribution with parameter $p=\mu\left(f(x_{1})-f(x_{2})\right)$).
In all our experiments, we use NN with $2$ hidden layers with width $50$, $\lambda=1.0$, $\delta=0.05$, $d=5$, $K=5$, and fixed value of $\nu_{T}=\nu=1.0$. For having a fair comparison, We choose the value of $\nu$ after doing a hyperparameter search over set $\{10.0,5.0,1.0,0.1,0.01,0.001,0.0\}$ for linear baselines, i.e., LinDB-UCB and LinDB-TS. As shown in Fig. 3, the average cumulative regret is minimum for $\nu=1.0$. Note that we did not perform any hyperparameter search for NDB-UCB and NDB-TS, whose performance can be further improved by doing the hyperparameter search.
As supported by the neural tangent kernel (NTK) theory, we can substitute the initial gradient $g(x;\theta_{0})$ for the original feature vector $x$ as $g(x;\theta_{0})$ represents the random Fourier features for the NTK [48]. In this paper, we use the feature vectors $g(x;\theta_{t})$ instead of $g(x;\theta_{0})$ and recompute all $g(x;\theta_{t})$ in each round for all past context-arm pairs. Additionally, compared to NTK theory, we have designed our algorithm to be more practical by adhering to the common practices in neural bandits [7, 8]. Specifically, in the loss function for training our NN (as defined in Eq. 1), we replaced the theoretical regularization parameter $\frac{1}{2}m\lambda\left\|\theta-\theta_{0}\right\|^{2}_{2}$ (where $m$ is the width of the NN) with the simpler $\lambda\left\|\theta\right\|^{2}_{2}$. Similarly, for the random features of the NTK, we replaced the theoretical $\frac{1}{\sqrt{m}}g(x;\theta_{t})$ with $g(x;\theta_{t})$.
Computational resources.
All the experiments are run on a server with AMD EPYC 7543 32-Core Processor, 256GB RAM, and 8 GeForce RTX 3080.
(a) Square (UCB)
(b) Square (TS)
(c) Cosine (UCB)
(d) Cosine (TS)
(a) Varying $d$ (Average)
(b) Varying $d$ (Weak)
(c) Varying $K$ (Average)
(d) Varying $K$ (Weak)
(a) $10(x^{\top}\theta)^{2}$ (Average)
(b) $10(x^{\top}\theta)^{2}$ (Weak)
(c) $20(x^{\top}\theta)^{2}$ (Average)
(d) $20(x^{\top}\theta)^{2}$ (Weak)
(e) $30(x^{\top}\theta)^{2}$ (Average)
(f) $30(x^{\top}\theta)^{2}$ (Weak)
(g) $40(x^{\top}\theta)^{2}$ (Average)
(h) $40(x^{\top}\theta)^{2}$ (Weak)
Figure 5: Comparisons of cumulative regret (average and weak) of different dueling bandits algorithms for non-linear reward functions.
(a) $cos(3x^{\top}\theta)$ (Average)
(b) $cos(3x^{\top}\theta)$ (Weak)
(c) $3cos(x^{\top}\theta)$ (Average)
(d) $3cos(x^{\top}\theta)$ (Weak)
(e) $5cos(x^{\top}\theta)$ (Average)
(f) $5cos(x^{\top}\theta)$ (Weak)
(g) $10cos(x^{\top}\theta)$ (Average)
(h) $10cos(x^{\top}\theta)$ (Weak)
Figure 6: Comparisons of cumulative regret (average and weak) of different dueling bandits algorithms for non-linear reward functions.
@@ Appendix D Leftover details from Section 6
Appendix E Broader Impacts
The contributions of our work are primarily theoretical. Therefore, we do not foresee any immediate negative societal impact in the short term. Regarding our longer-term impacts, as discussed in Section 4, our algorithms can be potentially adopted to improve online RLHF. On the positive side, our work can lead to better and more efficient alignment of LLMs through improved online RLHF, which could benefit society. On the other hand, the potential negative societal impacts arising from RLHF may also apply to our work. On the other hand, potential mitigation measures to prevent the misuse of RLHF would also help safeguard the potential misuse of our algorithms.

### v2

@@ D.2 Theoretical Analysis
In this section, we show the proof of 5 and 6 for neural contextual bandits with binary feedback (Appendix D). We can largely reuse the proof from Appendix A, and here we will only highlight the changes we need to make to the proof in Appendix A.
@@
To begin with, in our analysis here, we adopt the same requirement on the width of the NN specified in Eq. 8. First of all, 2 still holds in this case, which allows us to approximate the unknown utility function $f$ using a linear function. It is also easy to verify that 3 still holds. As a consequence, 4 and 5 both hold naturally.
D.2.1 Proof of Confidence Ellipsoid
@@
in which the inequality follows from 4. Then, the proof in Eq. 17 can be reused to show that
@@
Next, making use of the $1$-sub-sub-Gaussianity of the sequence of noise $\{\varepsilon_{s}\}$ and Theorem 1 from Abbasi-Yadkori et al. (2011), we can show that with probability of at least $1-\delta$,
@@
Denote $\varphi(x)=g(x;\theta_{0})$. Recall that 2 tells us that $f(x)=\langle g(x;\theta_{0}),\theta_{f}-\theta_{0}\rangle=\langle\varphi(x),\theta_{f}-\theta_{0}\rangle$ for all $x\in\mathcal{X}_{t},t\in[T]$. To begin with, for all $x\in\mathcal{X}_{t},t\in[T]$ we have that
@@
in which we have used 6 (reproduced in Eq. 40) in the last inequality. Now making use of the equation above and 5, we have that
@@
in which the last inequality follows from Eq. 41 and 5.
@@ D.2.2 Regret Analysis
D.3 Experiments for Neural GLM Bandits
@@
Appendix E Theoretical Insights for Reinforcement Learning with Human Feedback
Our algorithms and theoretical results can also provide insights on the celebrated reinforcement learning with human feedback (RLHF) algorithm (Chaudhari et al., 2024), which has been the most widely used method for the alignment of large language models (LLMs). In RLHF, we are given a dataset of user preferences, in which every data point consists of a prompt and a pair of responses generated by the LLM, as well as a binary observation indicating which response is preferred by the user. Following our notations in Section 2, the action $x_{t,1}$ (resp. $x_{t,2}$) corresponds to the concatenation of the prompt and the first response (resp. second response). Of note, RLHF is also based on the assumption that the user preference over a pair of responses is governed by the BTL model (Section 2). That is, the binary observation $y_{t}$ is sampled from a Bernoulli distribution, in which the probability that the first response is preferred over the second response is given by $\mathbb{P}\left\{x_{t,1}\succ x_{t,2}\right\}=\mu\left(f(x_{t,1})-f(x_{t,2})\right)$. Here $f$ is often referred to as the reward function, which is equivalent to the latent utility function $f$ in our setting (Section 2).
Typically, RLHF consists of two steps: (a) learning a reward model using the dataset of user preferences and (b) fine-tuning the LLM to maximize the learned reward model using reinforcement learning. In step (a), same as our algorithms, RLHF also uses an NN $h$ (which takes as input the embedding from a pre-trained LLM) to learn the reward model by minimizing the loss function (1). The accuracy of the learned reward model is crucial for the success of RLHF (Chaudhari et al., 2024). Importantly, our 1 provides a theoretical guarantee on the quality of the learned reward model $h$, i.e., an upper bound on the estimation error of the estimated reward differences between any pair of responses. Therefore, our 1 provides a theoretically principled measure of the accuracy of the learned reward model in RLHF, which can potentially be used to evaluate the quality of the learned reward model.
In addition, some recent works have proposed the paradigm of online/iterative RLHF (Bai et al., 2022; Menick et al., 2022; Mehta et al., 2023; Ji et al., 2024; Das et al., 2024) to further improve the alignment of LLMs. In online RLHF, the RLHF procedure is repeated multiple times. Specifically, after an LLM is fine-tuned to maximize the learned reward model, it is then used to generate pairs of responses to be used to query the user for preference feedback; then, the newly collected preference data is added to the preference dataset to be used to train a new reward model, which is again used to fine-tune the LLM. In this case, as the alignment of the LLM is improved after every round, the newly generated responses by the improved LLM are expected to achieve progressively higher reward values, which have been shown to lead to better alignment of LLMs (Bai et al., 2022; Menick et al., 2022; Mehta et al., 2023; Ji et al., 2024; Das et al., 2024). In every round, we can let the LLM generate a large number of responses (i.e., the actions in our setting, see Section 2), from which we can use our algorithms to select two responses $x_{t,1}$ and $x_{t,2}$ to be shown to the user for preference feedback. In addition, our algorithm can also potentially be used to select the prompts shown to the user, which correspond to the contexts in our problem setting (Section 2). Our theoretical results guarantee that our algorithms can help select responses with high reward values (2 and 3). Therefore, our algorithms can be used to improve the efficiency of online RLHF.
Appendix F Response Optimization in Large Language Models (LLMs)
As demonstrated in recent works (Bai et al., 2022; Menick et al., 2022; Mehta et al., 2023; Das et al., 2024; Ji et al., 2024), using responses with higher reward values can significantly improve the alignment of LLMs, especially in online/iterative RLHF where the responses generated by the initial LLM tend to have low rewards. Motivated by this, we use our proposed algorithms, NDB-UCB and NDB-TS, to select the responses with higher estimated reward values for a given prompt. In our experiment, we use the following setting from Lin et al. (2024): In each iteration, a user provides a prompt (context in our setting) to ChatGPT, and then ChatGPT generates $50$ responses (arms in our setting). We use Sentence-BER (Reimers & Gurevych, 2019) to get embedding representation for each prompt-response pair. Specifically, for iteration $t$, we use $x_{t,i}=(c_{t},a_{t,i})$ to denote the embedding representation generated by Sentence-BERT, where $c_{t}$ denotes the prompt and $a_{t,i}$ represents the $i$-th response. Of note, we adopt a reward model that is pre-trained using the Anthropic Helpfulness and Harmlessness datasets (Bai et al., 2022). For every prompt-response pair $x_{t,i}$, we use the output from this pre-trained reward model as the value of the unknown reward function $f(x_{t,i})$. This approach allows us to simulate the preference feedback in this experiment. Note that our approach to simulating preference feedback is common in the literature (Dwaracherla et al., 2024).
Figure 9: Scores of different algorithms for response optimization.
The embeddings $x_{t,i}$ are used as input to a trained neural network (NN) that estimates the latent reward model. Our proposed algorithms compute the reward estimate for each prompt-response pair to select the first response (using Eq. 2) and then use it with the optimistic reward estimates of other responses to select the second response (using either Eq. 3 or TS-based selection criterion for the second response). This way, our algorithms will ensure the selection of responses with consistently high reward estimates. As shown in Fig. 9, we compare the score (calculated as the sum of rewards for selected responses divided by the number of iterations) after $T$ iterations (multiple of 10) of NDB-UCB and NDB-TS and against two baselines: Linear variants (which uses a linear model to estimate latent reward model instead of an NN): Lin-UCB and Lin-TS, and Random (where two responses are selected randomly). The results clearly show that our algorithms achieve the highest score (especially NDB-UCB), demonstrating the superior performance of our proposed algorithm in identifying high-reward responses.
The experimental results shown in Fig. 9 corroborate our theoretical results, which guarantee the selection of responses with high reward values, as also discussed in Section 4.2 and Appendix E. This result further shows that our proposed algorithms can be used to improve the quality of the human preference dataset, which could then be used to improve the efficiency of online RLHF/DPO for LLM alignment. We leave it to future works to verify that the preference dataset with high-reward responses collected using our algorithms can indeed lead to better LLM alignment, which is beyond the scope of the current work (since the focus of this work is primarily theoretical) and will require significantly more extensive experiments.